30% of all sefdels have cele televison U.ร. In a random Sample of s as In a random LO 2
Answer #1

Here we have

n=15, p=0.70

Let X is a random variable shows the number of households that will have cable television. Here X has binomial distribution with parameters n=15 and p=0.70.

(i)

PlX-10) _ (10)(0.70)⑩(1-0.70)15-10-0.2061 ) (0.70)10(1-0.70)15-10-0.2061 = 10)

(ii)

15 / 15 P(X > 13) (0.70) 0.70)150.126s

(iii)

P(X < 13) 1-P(X > 13) 1-0.1268 = 0.8732

(iv)

15 P(10 < X < 12)-Σ ( ) (0.70)(1-0.70)15-1 = 0.5948 70) 15-1 P(10< X < 12)